OpenStudy (anonymous):

Some of the lowest pitches attainable on a musical instrument are achieved on the world’s largest pipe organs. What is the length of an organ pipe that is open on both ends and has a fundamental frequency of 8.75 Hz when the speed of sound in air is 341 m/s? (a) 9.83 m (c) 21.2 m (e) 32.4 m (b) 19.5 m (d) 29.3 m

OpenStudy (anonymous):

T = 1/f T - periof, f - frequency T = 0,114 sec L = V * T L = 38,97 m since it's open on both sides, then divide by 2 and the answer is 19,485 m So the answer is B.